### v1

# Beyond Recommendations: From Backward to Forward AI Support of Pilots’ Decision-Making Process

Abstract.
AI is anticipated to enhance human decision-making in high-stakes domains like aviation, but adoption is often hindered by challenges such as inappropriate reliance and poor alignment with users’ decision-making. Recent research suggests that a core underlying issue is the recommendation-centric design of many AI systems, i.e., they give end-to-end recommendations and ignore the rest of the decision-making process. Alternative support paradigms are rare, and it remains unclear how the few that do exist compare to recommendation-centric support. In this work, we aimed to empirically compare recommendation-centric support to an alternative paradigm, continuous support, in the context of diversions in aviation. We conducted a mixed-methods study with 32 professional pilots in a realistic setting. To ensure the quality of our study scenarios, we conducted a focus group with four additional pilots prior to the study. We found that continuous support can support pilots’ decision-making in a forward direction, allowing them to think more beyond the limits of the system and make faster decisions when combined with recommendations, though the forward support can be disrupted. Participants’ statements further suggest a shift in design goal away from providing recommendations, to supporting quick information gathering. Our results show ways to design more helpful and effective AI decision support that goes beyond end-to-end recommendations.
human-AI interaction, decision support tools, intelligent decision support, AI-assisted decision-making, human-AI decision-making, imperfect AI, decision support paradigms, continuous support, process-oriented support, aviation
journal: PACMHCIjournalvolume: 8journalnumber: CSCW2journalyear: 2024publicationmonth: 10copyright: rightsretainedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information systems Decision support systemsccs: Human-centered computing Interaction paradigms
Figure 1. Conceptual overview of the two decision support paradigms that we compare: recommendation-centric support and continuous support. While the former pushes pilots to reason backward from a decision recommendation, the aim of the latter is to support pilots in a forward direction. Instead of only popping up during an emergency, the system continuously helps pilots to evaluate their surroundings even during normal flight. The system does not give recommendations to avoid biasing pilots. But given continuous support, it may be possible to add recommendations in an emergency while still allowing pilots to reason forward, since pilots are already engaged with the system when they see the recommendations.
[…]
Schematic diagram of both support concepts.
1. Introduction
AI is projected to improve human decision-making in various high-stakes domains, such as healthcare (Yu et al., 2018), finance (Cao, 2022), or law enforcement (Raaijmakers, 2019). Another domain is aviation (EASA, 2023), where AI is expected to not only increase decision efficiency, but also safety, as faulty decision-making is one of the main reasons for accidents in aviation (Shappell et al., 2007), motivating research at the intersection between novel technologies and human factors (Feger et al., 2022). One type of decisions in aviation are diversion decisions. A diversion is when a flight is unable to reach its planned destination, e.g. due to a technical failure, a medical emergency, or adverse weather conditions. It is the pilots’ responsibility to decide on an alternate airport to divert to. While diversions are rare, they are very disruptive and costly for operations (Lewis et al., 2021). Poor diversion decisions can further increase the cost or even impact flight safety. Today, pilots have various tools to support them during diversions, but these are often cumbersome to use and not integrated with each other. Diversions are therefore one primary use case where pilots seek better support and where they can imagine AI assistance (Würfel et al., 2023).
Yet, in spite of good machine performance, real-world adoption of AI is often difficult (Schlögl et al., 2019; Varghese, 2020). In controlled studies, failure to achieve complementary performance is often observed (Bansal et al., 2021; Liu et al., 2021; Fok and Weld, 2023), i.e., the combination of human and AI performs worse than one of them alone. This is the result of inappropriate reliance, including both overreliance (human relies on AI even when it is disadvantageous to do so) (Bansal et al., 2021; Liu et al., 2021; Jacobs et al., 2021b; Schmidt and Biessmann, 2020; Bussone et al., 2015; Buçinca et al., 2021; Lai and Tan, 2019; Poursabzi-Sangdeh et al., 2021; Wang and Yin, 2021; Green and Chen, 2019) and underreliance (human rejects AI even when it would be beneficial to rely on it) (Dietvorst et al., 2015; Cheng and Chouldechova, 2023; Castelo et al., 2019; Prahl and Van Swol, 2017).
In practice, AI support often turns out less useful to decision makers than imagined (Yang et al., 2016; Blomberg et al., 2018; Kawakami et al., 2022a). A growing number of formative studies on real-world tasks with domain experts (Kawakami et al., 2022a, b; Jacobs et al., 2021a; Sivaraman et al., 2023; Bach et al., 2023; Burgess et al., 2023; Kaltenhauser et al., 2020; Yang et al., 2016; Zhang et al., 2023b) try to understand what hinders effective use of AI decision support. A frequent problem is that AI decision support is usually designed to be recommendation-centric, where the primary functionality of the system is to give end-to-end decision recommendations, i.e., the system suggests a possible end result straight from its input data. By directly jumping to the end result, these systems only support the very end of the decision-making (Buçinca et al., 2022), ignoring the entire process leading up to the decision (Yang et al., 2016; Zhang et al., 2023b, a).
While the limitations of recommendation-centric support are becoming increasingly more apparent, there is a notable lack of effective solutions to the identified problems. Current research is mostly limited to proposing alternatives for recommendation-centric support on a conceptual level (Zhang et al., 2021; Buçinca et al., 2022; Cabitza et al., 2021; Miller, 2023; Koon, 2022), but concrete examples are few and far between, with most examples stemming from healthcare (Cai et al., 2019; van Berkel et al., 2021; Yang et al., 2019; Lindvall et al., 2021; Gu et al., 2023; Zhang et al., 2023b). Even fewer works evaluate these alternative support paradigms in comparison to recommendation-centric decision support.
One of the few studies exploring alternative decision support roles for AI is Zhang et al.’s work on diversion assistance (Zhang et al., 2023a). Their key insight was that diversion decisions are not a point, but a process in which pilots take proactive actions. Even during normal flight, pilots constantly establish situation awareness (SA) and prepare a valid plan B, should an emergency occur. Continuously supporting these proactive actions via unobtrusive hints was found to be a promising support role for AI, but it remains unclear how effective it would be in practice, especially compared to the established recommendation-centric paradigm.
In this paper, we sought to empirically compare continuous against recommendation-centric support (see Figure 1) in terms of their effects on pilots’ decision-making process, decision outcomes, and decision time in a realistic task setting. We conducted a mixed-methods study with 32 professional pilots, where pilots made a series of diversion decisions with either recommendations, continuous support, a combination of both, or a baseline system with no AI. We aimed for higher ecological validity than in typical AI-assisted decision-making studies. To this end, we validated and refined our scenarios in a focus group with four additional pilots prior to the study.
Our results challenge the common assumption that AI decision support should be recommendation-centric. Continuous support allowed pilots to think more beyond the limits of the system, was better accepted by pilots, and led to faster decisions when combined with recommendations. Our paper makes the following three contributions:
  1. (1)
    We add to the rare examples of evaluative studies of AI-assisted decision-making with experts on a real-world task, in a domain that is understudied in the HCI community.
  2. (2)
    We conduct one of the first empirical comparisons between recommendation-centric and alternative forms of AI decision support, demonstrating the importance and potential of thinking beyond the typical recommendation-centric paradigm.
  3. (3)
    Based on our results, we propose a framework for process-oriented decision support as alternative to recommendation-centric support. We provide continuous support as well as further suggestions from our participants as concrete implementations of process-oriented support for diversions, but we consider the framework to be applicable in other domains beyond aviation as well.
2. Background and Related Work
We outline recent work on recommendation-centric decision support in Section 2.1 as well as alternatives to this dominant support paradigm in Section 2.2. We then describe the gap and the research questions we address in Section 2.3.
2.1. Recommendation-Centric Decision Support
The most common strategy to help decision makers work better with recommendation-centric AI is to add explanations of how the model works (Wang and Yin, 2021; Bussone et al., 2015; Bansal et al., 2021; Zhang et al., 2020; Lai and Tan, 2019) and other model information, such as model confidence (Zhang et al., 2020; McGuirl and Sarter, 2006; Rechkemmer and Yin, 2022) or the stated model accuracy (He et al., 2023; Rechkemmer and Yin, 2022). The goal is to help people to rely appropriately on AI recommendations (Schemmer et al., 2023) by giving cues about when they may be beneficial or detrimental to rely on. Results have been mixed so far, as especially explanations are prone to induce blind trust (Bansal et al., 2021; Eiband et al., 2019) and hence overreliance, even among domain experts (Bussone et al., 2015; Jacobs et al., 2021b). Recently, Vasconcelos et al. (Vasconcelos et al., 2023) have shown that explanations can reduce overreliance under certain conditions, but it is questionable how often these conditions are valid in real applications (Fok and Weld, 2023). Communicating model confidence appears more promising, as it has repeatedly been shown to improve appropriate reliance (Zhang et al., 2020; McGuirl and Sarter, 2006; Bansal et al., 2021; Prabhudesai et al., 2023); but this relies on well-calibrated confidence scores, which are often difficult to achieve—models can be wrong with high confidence.
The limited success of adding model information appears to be due to people not engaging cognitively with it (Buçinca et al., 2021; Gajos and Mamykina, 2022). One way to increase engagement is to employ cognitive forcing interventions, such as showing recommendations only after users made an initial decision (Buçinca et al., 2021; Fogliato et al., 2022), introducing a waiting time until recommendations are shown (Buçinca et al., 2021; Park et al., 2019), or forcing users to wait before they can proceed to the next task (Rastogi et al., 2022). While effective, these interventions negatively impact user experience (Buçinca et al., 2021; Fogliato et al., 2022; Park et al., 2019). Liu et al (Liu et al., 2021) explored the use of interactive explanations to increase engagement, though it did not reduce overreliance in their case.
Recently, an increasing number of voices call the entire premise of recommendation-centric support into question. Koon (Koon, 2022) emphasizes that decisions are often complex, and that condensing them into an AI recommendation is necessarily reductionist. At the same time, since recommendations are hard to appropriate, users often struggle to combine them with the wider context knowledge they have (Cabitza et al., 2021). From a cognitive science perspective, Miller (Miller, 2023) argues that recommendation-centric support does not align with the cognitive processes of human decision-making. Instead, recommendations take control away from decision makers. Similarly, Wang et al. (Wang et al., 2019) and Zhang et al. (Zhang et al., 2021) caution against error-prone backward reasoning from the end result back to the input data, which is facilitated by a fixation on end-to-end recommendations. All of these authors call for alternative approaches to AI decision support that are less centered on recommendations.
2.2. Alternative Forms of AI Decision Support
One stream of work that de-emphasizes recommendations rethinks the purpose of explanations. Instead of explaining the AI model, explanations can provide information that is of natural interest in a decision, such as domain-specific information (Yang et al., 2023; Lim et al., 2023), or the socio-organizational context of a decision (Ehsan et al., 2021). These explanations situate recommendations within the primary decision-making task, rather than diverting attention to the secondary task of understanding the AI model.
Other authors propose entirely different roles for AI than providing end-to-end recommendations. Alternative AI support paradigms are far from new (Woods, 1986), but have been largely ignored by current research. This is arguably more due to technical feasibility rather than consideration for human needs, since with modern AI methods, it is straightforward to formulate many decision tasks as end-to-end predictions (Buçinca et al., 2022; Shneiderman, 2020). As for alternatives to providing recommendations, on a conceptual level, Cabitza et al. (Cabitza et al., 2021) propose to frame AI as “knowledge artifact functions” which support people in their collaborative decision-making. Zhang et al. (Zhang et al., 2021) put forward what they call “forward-reasoning decision support”, where users form decisions themselves, augmented by rich interactions with AI tools. In a similar, but more concrete way, Miller (Miller, 2023) proposes the concept of “evaluative AI”, which focuses on helping decision makers to evaluate different hypotheses.
In essence, all of these proposals aim to help people to make decisions through forward reasoning, allowing people to start from the context at hand and to use their domain expertise to reach a decision. This is in contrast to recommendation-centric support, which pushes people to reason backward from the recommendation. In fact, cognitive forcing can be seen as a way to encourage forward reasoning as well by pushing people to think independently from AI recommendations. However, with cognitive forcing interventions, people get no support while making their independent decisions, which reduces the supportive value of the AI (Buçinca et al., 2021; Park et al., 2019; Fogliato et al., 2022). The above concepts aim to facilitate forward reasoning while supporting users’ decision-making processes.
Beyond abstract concepts, concrete examples for supporting decisions without relying on end-to-end recommendations are often found in healthcare. Lindvall et al. (Lindvall et al., 2021) designed a system for tumor assessment that navigates pathologists to potentially tumorous image regions for them to review, without revealing whether or how confidently the model classifies the pixels as tumor. Crucially, the classification threshold was not set to optimize accuracy, but sensitivity. Consequently, if pathologists did not find a tumor in any of the suggested regions, they could be relatively sure that the rest of the image also contains no cancer. Zhang et al. (Zhang et al., 2023b) studied the case of sepsis diagnosis, where an existing AI tool only addresses the final stage of the decision-making by offering a sepsis risk score and sending alerts above a certain threshold. The authors proposed a redesign where the system suggests lab tests that would help to reduce uncertainty about patients’ future conditions.
In the context of aviation, Zhang et al. explored how to support diversions decisions (Zhang et al., 2023a). Their system does not recommend airports, but continuously provides unobtrusive local hints about potential limitations at the surrounding airports. The system provides this support of pilots’ SA also during normal flight when there is no sign of an emergency yet.
2.3. Summary and Research Questions
There is a remarkable difference between the studies mentioned in the previous two sections: The studies on recommendation-centric support in Section 2.1 were almost all built on simple, artificial tasks with lay persons as participants. Such studies make up the majority of research in AI-assisted decision-making (Lai et al., 2023). In contrast, the alternative support paradigms in Section 2.2 all stemmed from studying complex real-world decisions with domain experts. This indicates a potentially significant gap between what is studied in large-scale controlled experiments and what is actually required by experts in real applications. Our study is situated right in this gap by conducting a controlled comparison between alternative decision support paradigms with domain experts.
The goal of our work is to empirically compare continuous support with typical recommendation-centric support. We further add a combination of both to the comparison, where the system provides continuous support during normal flight, but gives recommendations in an emergency. Figure 1 contrasts how the two paradigms conceptually fit into pilots’ decision-making, based on the FOR-DEC model (Hörmann, 1994). FOR-DEC is a prescriptive model used by many airlines to train pilots to make decisions in a structured manner. It is an acronym for the steps to follow during decision-making: facts, options, risks & benefits, decision, execution, check. The dash in the middle signals a pause for reflection before making the decision. Figure 1 only covers the first four steps of the model, since execution and check are beyond the scope of our work.
The focus of our empirical comparison is to assess whether pilots reason differently with continuous support and recommendation-centric support, and how this affects pilots’ perceptions about the system as well as decision outcomes, namely overreliance and decision time. As outlined in Section 2.1, the former is one of the major concerns in recommendation-centric support. The latter is of interest since aviation is highly dynamic, which limits the time available for decision-making. We therefore pose the following research questions:
  • RQ1: How do pilots integrate the different support paradigms into their workflow?
  • RQ2: How do the support paradigms differ in terms of overreliance?
  • RQ3: How do the support paradigms differ in terms of decision time?
  • RQ4: How do pilots perceive the different support paradigms?
We hypothesize that recommendations induce backward reasoning, while continuous support facilitates forward reasoning. We further hypothesize that combining both allows to give recommendations while still enabling forward reasoning. Hence, we expect continuous support to lead to less overreliance than recommendations, but at the expense of longer decision times. We expect the combination of both to combine their advantages, i.e., less overreliance and shorter decision times.
Most closely related to our work is a study by Smith et al. (Smith et al., 1997), who compared the effects of three different versions of a flight planning tool on users’ performance. Apart from being a different task, the major difference to our work is that their system versions represent consecutively increasing levels of automation, while we study alternative support paradigms.
3. Diversion Assistance System
In this section, we present the design of the four versions our diversion assistance system (Section 3.1), followed by the apparatus used to conduct our study (Section 3.2).
3.1. System Design and Variants
[…]
Table with four versions of the diversion assistance.
[…]
Screenshots of the system.
[…]
3.2. Apparatus
As simulation environment, we chose X-Plane 11222https://www.x-plane.com/, running the Airbus A320 Ultimate aircraft model by Flight Factor [^3] https://flightfactor.aero/. We supplemented the simulator with hardware controls, including a sidestick and a throttle quadrant from Thrustmaster [^4] https://www.thrustmaster.com/, as well as an MCDU [^5] Multipurpose Control and Display Unit, the input device for the flight management system and other computer systems. and an FCU [^6] Flight Control Unit, the control panel used to control the autopilot. from Skalarki [^7] https://www.skalarki-electronics.com/, to reduce reliance on a mouse for interacting with the aircraft. The DAS was implemented as an interactive mockup in the prototyping tool Framer [^8] https://www.framer.com/, with custom components written in React. It runs on a second-generation Surface Pro tablet next to the simulator computer, mimicking an EFB [^9] Electronic Flight Bag, a portable device to store and display flight-relevant data and to run assistive applications.. The mockup communicates via Websocket with a Python back end on the simulator computer, which acts as an interface to X-Plane to read out live flight data. Weather and ground facilities at each airport are hard-coded for each scenario. The Python back end further triggers the emergencies of the study scenarios in X-Plane based on timers. The entire setup is shown in Figure 4. Note that despite the remarkable realism of the elements we used in our simulation setup, some notable deviations from real flights remained. We describe these deviations and discuss their implications in Section 6.4.
Figure 4. Overview of the study apparatus.
[…]
Schematic overview of study apparatus.
The AI recommendations are generated with a manually-tuned scoring function that assigns each airport a score, calculated as a weighted sum of subscores for each criterion. Each subscore is calculated using a piecewise linear function that penalizes unfulfilled criteria more heavily than it rewards overfulfilled criteria. Using this scoring function, the recommendations react dynamically to what is happening in the simulator and to the criteria the pilot defines. While it is not machine learning, given the large number of criteria, it is still hard to comprehend what exactly causes a certain airport to be recommended or not, just as with an opaque deep neural network.
4. Methods
Besides the DAS, valid and informative scenarios were the key component of our study. To ensure the quality of our scenarios, we first conducted a focus group to discuss and flesh out our initial scenario outlines with professional pilots. We subsequently tested our scenarios and system design with pilots before finally running our main study. In this section, we describe these steps and our methodology in detail, which are shown in Figure 5.
Figure 5. Overview of the study parts.
[…]
Schematic overview of the study parts.
4.1. Focus Group and Study Scenarios
There is no inherent right or wrong in most diversion decisions, so one goal of the focus group was to obtain a reference for how pilots would decide in each scenario and for which reasons. The other goal was to understand how to design the details of the scenarios to fit our intentions. We recruited four professional pilots for the focus group (one captain, three first officers; all male; median age: 29.5 years (IQR 28.5–34.75); median flight hours: 1000 hours (IQR 392.5–3750); details in Appendix B, Table 3). All four work at the same airline, but have past experience in four additional airlines. We showed them the initial outline of each scenario and asked them to discuss how they would decide. The focus group had a duration of 90 minutes, and participants were compensated with 150 EUR ($\approx$ 164 USD) each, which is a typical rate for professional pilots, given the difficulty to recruit them.
After the focus group, we adjusted the scenario outlines and filled in the details according to the insights gained. In total, we designed three scenarios, with the aim to cover a good range of common reasons for diversions as well as different potential failure modes of the DAS, as recommended by Roth et al. (Roth et al., 2022). In particular, we aimed to construct one scenario each where the DAS performs well (Scenario 1), suggests an airport that pilots tend to disagree with (Scenario 2), and suggests a solution that is subpar, but for a reason that is not immediately obvious (Scenario 3).
4.1.1. Scenario 1: Single engine failure
Figure 6. Single engine failure scenario. The flight is from Geneva to London. Figure based on SkyVector [^10] https://skyvector.com/.
[…]
Screenshot of single engine failure scenario.
In the first scenario, a flight from Geneva, Switzerland to London Heathrow suffers a single engine failure during climb (footnote 10). The scenario is based on incident reports found with the CAROL query tool of the National Transportation Safety Board [^11] https://carol.ntsb.gov/. The focus group participants pointed out that the terrain around Geneva may be challenging with an engine failure, but returning to Geneva is likely preferable for the company. Charles de Gaulle on the other hand is a good option since it is a major hub with excellent infrastructure.
Our intention for this scenario was that the DAS would recommend Charles de Gaulle as an option that pilots would agree with. Beginning with this scenario was meant to establish pilots’ initial trust into the system, since in reality, such a system would perform well in most situations. To make the decision less ambiguous and hence the recommendation more acceptable, we added a moderately high crosswind component to Geneva. In the scenario, we triggered an engine failure and displayed a popup message within X-Plane, reading “Engine 1 has failed. No engine relight [^12] In reality, pilots would try to relight the engine. We asked them to skip this to simplify the scenario, as it would not have added any value to our study.. Please secure the engine and proceed with the diversion decision.” Pilots were asked to handle the engine failure within X-Plane as they would in reality to keep them in their known workflow.
4.1.2. Scenario 2: Passenger medical emergency
Figure 7. Medical emergency scenario. The flight is from Tenerife to Munich. Figure based on SkyVector.
[…]
Screenshot of passenger medical emergency scenario.
The second scenario was a flight from Tenerife, Spain to Munich, Germany, with a passenger having a heart attack while the flight is above the sea in the middle between Tenerife and the European continent (Figure 7). The emergency was announced through a popup message in X-Plane: “The passenger on 11C has vomited, complains of extreme chest pains, is pale and sweating. This has been going on for a few minutes now. A doctor (general practitioner) sitting next to the passenger reacted immediately. He suspects a heart attack.” This scenario was based on a route that is notorious among short-haul flight pilots due to the long flight over open sea, with much fewer diversion options than usual on European short-haul flights.
Our expectation was that Casablanca, Morocco would be the obvious option since it is the closest, making the decision too easy. Hence, we thought about adding a reason to make Casablanca unattractive, like a nation-wide healthcare strike. However, our focus group participants said they would either fly to Faro, Portugal, or return to Tenerife, depending on which is faster, since they would prefer to stay within Europe. This would make it easier to continue the flight and was also assumed to allow for better medical treatment for the passenger. We therefore decided not to add any factor to make Casablanca unattractive. Still, the DAS would recommend Casablanca to probe for pilots’ reactions when the system recommends an option that they are not comfortable with.
4.1.3. Scenario 3: Airport closure
Figure 8. Airport closure scenario. The flight is from Sharm El Sheikh to Hamburg. Figure based on SkyVector.
[…]
Screenshot of airport closure scenario.
This last scenario was based on the personal experience of one participant in Zhang et al.’s study (Zhang et al., 2023a). The flight was from Sharm El Sheikh, Egypt, to Hamburg, Germany (Figure 8). About an hour before landing, participants were given the information that the destination had closed down, again via a popup message: “Due to a power outage, EDDH [^13] The ICAO (International Civil Aviation Organization) airport code for Hamburg. had to close completely. It will remain closed for at least the next few hours.”
The intention for this scenario was that the DAS would recommend an option that is good at first glance, but problematic at second thought, in order to probe for how much participants would over-rely. The system would recommend Hanover, Germany, which is the typical alternate for Hamburg. It also has good ground connections to Hamburg for the passengers and is therefore a plausible recommendation. However, the airport closure affects the entire air traffic coming into Hamburg. Many planes would try to divert to Hanover, which in the real event had caused Hanover to run full. The DAS does not consider this traffic factor.
Unprompted, focus group participants discussed the traffic situation in Hanover as an important consideration, confirming that pilots would realistically think about it in such a situation. Participants said that instead of Hanover, all the en-route airports like Berlin, Dresden, or Leipzig would be valid options. We added slightly unfavorable, but workable weather conditions at these airports, to make the system recommendation Hanover seem more attractive at first glance.
4.2. Main Study
The main study had a between-subject design where each participant was assigned to one of the four DAS variants in Figure 2 and completed all three scenarios in the order of Section 4.1. At the beginning of the study, pilots got a detailed introduction into X-Plane and the DAS. Before starting each scenario, pilots were shown maps similar to those in footnote 10Figure 8, with the flight plan, the departure and destination airport, as well as the position of the aircraft at the beginning of the scenario. The scenarios started a couple minutes before the incident to allow participants to familiarize themselves with the situation. Participants were allowed to freely interact with the flight simulation, but were asked to only call up the DAS (with Recommendations and Baseline) or enter emergency mode (with Continuous Support and Recommendations + Continuous Support) when the incident happened. With Continuous Support and Recommendations + Continuous Support, pilots were allowed to freely interact with the normal flight mode before the incident.
Participants were asked to think aloud and clearly announce which airport they would divert to; they did not need to execute the diversion. The pilots were encouraged to think beyond what they saw in the DAS and consider whatever they would in a real flight, including which other stakeholders they would contact, like their company or air traffic control. They were further told that whatever is not shown in the table is not considered by the AI.
[…]
We pilot-tested the study procedure with two of the focus group participants from Section 4.1 (see Appendix B, Table 3), whom we also paid 150 EUR each. Following the pilot test, we made small adjustments to the DAS user interface and to the third scenario. Most importantly, both pilot testers asked for preferences of the company in Scenario 3. We therefore decided that if participants would ask for this information, we would tell them that the airline would prefer a diversion to either Hanover, Berlin, or Düsseldorf. Hanover was the top recommendation of the AI, Düsseldorf was the third recommendation, and Berlin was not recommended due to slightly worse weather.
4.3. Data Analysis
We took participants’ decisions and the time they needed for their decisions as outcome measures. As decision time, we took the time between calling up the system or entering emergency mode and announcing the decision. We further noted the reasons for the decisions from the think-aloud protocols to cover both outcome and process measures, as recommended by Roth et al. (Roth et al., 2022).
To analyze the qualitative data, we transcribed the think-aloud protocols and exit interviews and coded both parts through thematic analysis (Blandford et al., 2016). An initial round of open coding was conducted independently by two authors on four transcripts, one per study condition. The two authors then discussed their initial set of 269 low-level codes and consolidated them into 18 code groups. The first author coded the rest of the transcripts, extending and revising the initial coding scheme when necessary. All additional and revised low-level codes in this step continued fitting into the 18 code groups. Finally, we identified three themes that reflect the code groups.
Additionally, we extracted usage patterns from the think-aloud protocols by reviewing the codes in their temporal order in each of the 96 decision instances (32 participants $\times$ 3 scenarios), referring back to the screen recordings where necessary for more context.
5. Results
For our main study, we recruited 32 professional pilots through snowball sampling (two captains, rest first officers; two female, rest male; median age: 31 years (IQR 30–35); median flight hours: 2500 hours (IQR 2000–4000); median of 1.5 self-performed diversions (IQR 0–3); details in Appendix B, Table 4). The pilots work in four German airlines, with past experience in eight additional airlines.
[…]
5.1. RQ1: Workflow Integration
We first describe the usage patterns we identified, followed by a comparison of how often they occurred across scenarios and system variants.
5.1.1. Usage patterns
All participants intuitively integrated the DAS into the FOR-DEC framework they are familiar with. We therefore distinguish the identified usage patterns on two overarching levels: the options considered by participants, and the strategies they used to decide between the options. The first directly mirrors the options step in FOR-DEC, while the latter is a combination of the risks & benefits and the decision steps, which were not always clearly distinguishable in the think-aloud protocols. We identified three distinct patterns for the options considered:
  • O1—First few options: Participants considered the first few options in the table. For Rec and Rec+Cont, these were the AI recommendations. For Cont and Baseline, these were the closest airports, since the table was sorted by time per default.
  • O2—First option: Like the above, but participants only considered the very first instead of the first few options in the table.
  • O3—Self-generated options: Participants generated their options independently from the order in the table, e.g. by sorting the table according to a column of interest, by looking for familiar airports, or by asking for company preferences.
We further identified twelve strategies that participants used to decide between the options they considered. The first five are general strategies that could be used across all variants:
  • S1—Narrow down: Participants carefully compared the considered options, ruling out options one by one until one remains.
  • S2—Recognize best: Participants immediately recognized the best among the considered options and focused on it without detailed comparison against the other options.
  • S3—Check if first option works: In case participants considered only the first option, they checked whether there was anything against it.
  • S4—Confirm first option is best: In case participants considered only the first option, they did a quick cross check to confirm that it was indeed the best, e.g. by glancing over the time to destination to see if the first option was significantly closer than the next ones.
  • S5—Check one after another: Participants checked whether there was anything against the first option. If yes, they checked the second one, and so on.
The seven remaining strategies capture how participants used the AI support elements available to them. The normal flight mode of the Cont and Rec+Cont variants was used in two different ways:
  • S6—Use prepared plan: Participants prepared a plan for a hypothetical emergency during normal flight. Even though they did not know what would happen, they could prepare on the level of e.g., “If something very urgent would happen, $x$ would be a good option.” When the emergency happened, participants used this plan to quickly reach a decision.
  • S7—Refine situation awareness: During normal flight, participants were aware of the general situation, such as the weather around them. When the emergency happened, participants did not review this information again, but only supplemented it with situation-specific information like the distance to the next hospital in Scenario 2.
The color highlights, which served as transparency for recommendations and as local hints for continuous support, were also used in two ways:
  • S8—Look for options without highlights: Participants reviewed highlights for their relevance, but gravitated toward options without highlights. This strategy was used for choosing among the considered options as well as for deciding which options to consider in the first place. For instance, participants looked at time to destination and quickly excluded those options with red highlights, which indicated that these airports were too far away.
  • S9—No highlights as confirmation: Participants had a decision in mind and felt confirmed if that option had no highlights. The difference to above is that there, the highlights were used as cue to guide the decision-making, while here, participants considered the highlights as a second opinion to their own independent reasoning.
Lastly, participants employed three distinct strategies for using recommendations:
  • S10—Recommendation as confirmation: Participants had a decision in mind and felt confirmed if that option was also recommended by the DAS. Some pilots using this strategy explicitly ignored the recommendations initially to reach a decision on their own first.
  • S11—Recommendation as fallback: Participants first checked their self-generated options. When they were not satisfied with any of them, they chose the AI-recommended option.
  • S12—Negotiate: Participants edited the criteria to see how it affected the recommendations. This strategy was triggered by one of two reasons, or a combination of both: Some pilots disagreed with the pre-defined criteria, while others were surprised that their favored option was not recommended. In the latter case, they tried to align the recommendations with their own opinion by removing or relaxing criteria they deemed uncritical. Those that successfully aligned the recommendations to their favored option took this as confirmation. One pilot did not succeed, which triggered further considerations in his decision-making.
5.1.2. Differences between scenarios and support paradigms
Figure 9. Usage patterns across conditions and scenarios. The counts exceed the number of participants because individual usage patterns are not mutually exclusive and participants typically employed more than one pattern. Grayed-out fields mean that the patterns are not applicable to the respective system variant.
[…]
A table of the different usage patterns.
Figure 9 gives an overview of the occurrences of each usage pattern across conditions and scenarios. Note that the real occurrences could be higher, since participants might have used certain strategies without any revealing verbalizations. Especially some of the AI usage strategies were more subconscious and therefore less likely to be verbalized.
The most obvious difference is between the options considered in different scenarios, where participants tended to consider the first few options (O1) in Scenario 1 and only the first option (O2) in Scenario 2, while they based their decision more heavily on self-generated options (O3) in Scenario 3. The difference was mostly due to the varying time criticality, as several participants explained during the exit interviews. The engine failure in Scenario 1 was of medium time criticality, the passenger emergency in Scenario 2 was extremely time-critical, while the airport closure in Scenario 3 was not time-critical at all.
Besides these scenario differences, there are also differences between the DAS variants. These are most salient in Scenario 2, where participants in the Cont and Rec+Cont groups benefitted the most from the possibility of detailed pre-planning with the normal flight mode. As a result, pilots in these groups could use their prepared plan (S6) or use recommendations as confirmation (S10) for their plan, indicating forward reasoning by these participants. Note that pilots using the Baseline and Rec variants also tried to establish SA during normal flight, as they would in reality, but could only rely on the much more limited possibilities offered by X-Plane. Consequently, when the emergency happened, pilots in the Rec group were more likely to take the system recommendation as starting point for review (S3, S4), which indicates backward reasoning. In the absence of AI support that helps to identify the best option, participants using the Baseline variant had a stronger tendency to review the first few options (O1) rather than only the first option (O2).
In Scenarios 1 and 3, the benefits of pre-planning during normal flight in Cont and Rec+Cont are less apparent. This can be explained by a disruption between normal flight and emergency decision-making in both scenarios. In Scenario 1, the disruption happened because pilots first handled the engine failure before entering the diversion decision, a procedure that took several minutes during which the position of the aircraft changed significantly. After securing the engine, pilots therefore had to slightly re-orient and could not directly use their SA from the normal flight. In Scenario 3, the disruption occurred since pilots mostly prepared for a time-critical emergency requiring a quick landing. What happened instead was a situation where it was more important to find a suitable airport near the destination so that participants’ preparations were not applicable.
Nevertheless, a difference is still observable between Rec and Cont in Scenario 1, indicating that continuous support encourages forward reasoning, while recommendations prompt backward reasoning. All participants using the Rec variant took the system recommendations as starting point (O1), while participants in the Cont group tended to consider self-generated options (O3), suggesting that they could build on their normal flight SA to some extent, despite the disruption of handling the engine failure. Furthermore, some participants in the Rec+Cont group limited their options to the system’s top recommendation (O2), which could have been because it conformed to their impressions of the available options from the normal flight phase.
In Scenario 3, a stark contrast is apparent between Baseline and Cont on the one hand—which did not have AI recommendations—and Rec and Rec+Cont on the other hand. Almost all participants in this scenario considered self-generated options (O3), mostly by asking for company preferences and by looking for familiar airports. In addition, those participants who had recommendations available to them also mostly considered them (O1), indicating a strong influence of the recommendations. However, from the think-aloud protocols, we were not able to tell whether they influenced participants’ reasoning differently between Rec and Rec+Cont, given that the observable recommendations usage patterns are very similar (S10, S11, S12). Still, the observed usage patterns presented in Figure 9 suggest that our hypothesis holds true according to which recommendations prompt backward reasoning, while continuous support facilitates forward reasoning, even though the latter can be derailed by disruptions between normal flight and emergency decision-making.
5.2. RQ2: Overreliance
Decision outcomes were quite uniform for the first two scenarios. For Scenario 1, this was to be expected, since the scenario was designed to be rather unambiguous for the reason given in Section 4.1.1. All participants except for C5 (Lyon) and B8 (Paris Orly) decided for Paris Charles de Gaulle in Scenario 1, which was also the system’s top recommendation. However, in Scenario 2, all participants except for B5 (Faro) chose Casablanca, which ran contrary to our expectation given the discussion of our focus group participants, as described in Section 4.1.2. Some participants in the main study, especially those who had flown in the region themselves before, emphasized that the bias against diverting to Africa is common among their colleagues, but unwarranted in this case. 14 participants across all conditions did mention that they would generally prefer to divert to Faro for the reasons also discussed by the focus group participants. In the end, the fact that the emergency was urgent and that Casablanca was around ten minutes closer tipped the scales for all of these pilots. The AI elements—which favored Casablanca—did not seem to have an effect on this decision, given that almost all participants in the Baseline group also decided for Casablanca.
As intended, Scenario 3 was less clear for participants, as shown in Table 1. Of most interest in terms of overreliance was how many participants decided for Hanover without considering the traffic situation, as explained in Section 4.1.3. While some participants did rule out Hanover themselves because of the traffic, other pilots either said they would ask air traffic control about the traffic density, or they considered traffic but expected it to be not too dense to land there. We told these participants that Hanover was already running out of capacity, as air traffic control would do in reality. We did so since we were not interested in how pilots would assess the traffic situation, but only whether they would consider it at all. The DAS does not include this information, so we were interested in whether this would lure pilots into overlooking this factor, or whether they would think beyond the limits of the system.
Table 1. Diversion decisions across conditions in Scenario 3. Hanover was to be avoided due to heavy traffic, which was not considered by the system.
Baseline Rec Cont Rec+Cont
To be avoided Hanover AI recommendation & company preference 1/8 6/8 4/8 3/8
Other Berlin company preference 7/8 - 4/8 3/8
Düsseldorf company preference - 1/8 - 1/8
Cologne - - - 1/8
Frankfurt - 1/8 - -
Figure 10 shows how the three AI system variants affected the probability for the overreliant behavior of choosing Hanover without considering traffic, as compared to the Baseline variant. Since we assumed that adding AI would increase the probability for overreliance, we performed one-sided Fisher’s exact tests to compare each AI variant with Baseline. We further report relative risks (RR) compared to Baseline with Wald normal approximation confidence intervals as effect sizes. Consistent with our hypothesis, there was a statistically significant increase of overreliance probability for Rec ($RR=6$, 95% CI [0.92, 39.18], $p=0.02<0.05$). For Cont ($RR=4$, 95% CI [0.56, 28.40], $p=0.14$) and Rec+Cont ($RR=3$, 95% CI [0.39, 23.07], $p=0.28$), the increase was smaller and not statistically significant, though these latter results have to be interpreted carefully given the small sample sizes and the resulting large confidence intervals. However, it appears like even without recommendations, pilots using the Cont variant could still be biased by the hints: Some airports in Scenario 3, including Berlin, had a braking action of “medium to good”, i.e., slightly wet runways, which the system highlighted in red. With the Baseline variant, which did not have highlights, participants acknowledged the wet runways with a short comment that it was uncritical; some even did not verbalize any thoughts about it at all. By contrast, with the AI variants, many participants found the red highlights hard to ignore, even though they deemed the braking action uncritical: “To me, red is always intuitively, it doesn’t look so good. That’s why I looked a bit at Hanover, even though ‘medium to good’ is fine.” (C8). This behavior is also reflected by the usage patterns in Figure 9, where some Cont participants looked for options without highlights (S8).
Figure 10. Probability of overreliance across conditions, where choosing Hanover in Scenario 3 without considering traffic conditions is counted as overreliance. Error bars denote 95% confidence intervals, estimated using the Wilson score interval method. (*) denotes $p<0.05$.
[…]
Point plot of overreliance occurrences.
In addition to the decisions, we also noted the reasons behind them. Table 2 gives an overview of the reasons why non-overreliant participants decided against Hanover, showing that traffic density was indeed the reason for most of them. Some participants decided against Hanover but not due to traffic. These pilots—all of whom saw no recommendations—favored Berlin for other reasons and focused on it without discussing Hanover in depth. Interestingly, the pilots in the Rec group who thought of traffic were two of the three pilots in that group who ignored the recommendations (i.e., not used the pattern O1 in Figure 9) to make their own independent thoughts. This suggests that forward reasoning was important for pilots to be able to think beyond the limits of the system.
Table 2. Reasons for not diverting to Hanover in Scenario 3. Participants could give more than one reason.
Baseline Rec Cont Rec+Cont
Traffic in Hanover 5/7 2/2 3/4 5/5
Berlin is closer to current location 2/7 - 1/4 -
Berlin has better connections for crew to return home - - 1/4 -
5.3. RQ3: Decision Time
The differences in usage patterns discussed in Section 5.1.2 also resulted in different patterns in decision times, as shown in Figure 11. We therefore analyzed each scenario separately through an analysis of variance (ANOVA). We further used Tukey’s HSD test for post-hoc tests and report both mean differences in seconds and Cohen’s $d$ as effect sizes.
(a) Scenario 1 (engine failure)
(b) Scenario 2 (passenger emergency)
(c) Scenario 3 (airport closure)
Figure 11. Decision time across conditions for all three scenarios. Error bars represent 95% confidence intervals. (+) denotes $p<0.1$, (*) denotes $p<0.05$, and (**) denotes $p<0.01$.
In Scenario 1, we found statistically significant differences in decision times at the $p<0.01$ level ($F(3,28)=5.15,p=0.006$), with a large effect size of $\omega^{2}=0.28$. Post-hoc tests revealed that all three AI system variants led to significantly faster decisions than Baseline: Rec was significantly faster than Baseline at the $p<0.1$ level ($\Delta=84.88$ s, 95% CI [-4.30 s, 174.05 s], $d=1.30$, 95% CI [-0.20, 2.80], $p=0.066$), Cont at the $p<0.05$ level ($\Delta=99.00$ s, 95% CI [9.83 s, 188.17 s], $d=1.52$, 95% CI [-0.02, 2.05], $p=0.025$), and Rec+Cont at the $p<0.01$ level ($\Delta=119.13$ s, 95% CI [29.95 s, 208.30 s], $d=1.82$, 95% CI [0.25, 3.40], $p=0.006$). The effect sizes can be considered large in all three comparisons. The differences between the AI variants were not statistically significant.
In Scenario 2, decision times differed significantly at the $p<0.1$ level ($F(3,28)=2.66,p=0.067$) with a medium effect size of $\omega^{2}=0.135$. Post-hoc test results showed that Rec+Cont was significantly faster than Rec at the $p<0.1$ level with a large effect size ($\Delta=82.00$ s, 95% CI [-5.61 s, 169.61 s], $d=1.28$, 95% CI [-0.22, 2.78], $p=0.073$). Other post-hoc comparisons were not statistically significant. The most notable difference to Scenario 1 was that the Baseline variant was similarly fast as Rec and Cont. This effect can be explained by the difference in time criticality, as discussed in Section 5.1.2. In Scenario 1, pilots compared several options in greater detail, which was facilitated by AI support; whereas in Scenario 2, they tended to only check the nearest airport to speed up the decision. This short check could be performed quickly even without AI support.
There were no statistically significant differences in decision times in Scenario 3 ($F(3,28)=0.544,p=0.66)$, with an effect size close to zero. This again can be explained by the factor of time criticality. Participants had no time pressure in Scenario 3 and hence took their time for in-depth comparisons, irrespective of which system variant they used.
Overall, the most notable observation was that Rec was surprisingly not faster than Cont; it even tended to be slightly slower in the first two scenarios, though not significantly. Interestingly, the combination Rec+Cont did lead to a statistically significant speedup in Scenario 2. This is consistent with the usage patterns in Figure 9: Many participants in this group already prepared during normal flight for Casablanca in case of a time-critical emergency. When the emergency happened, the system confirmed their prepared plan (S10, forward reasoning), reducing the time to verify the plan. With Rec, participants got the same recommendation, but since they were less prepared, they had to verify the recommendation first (S3, S4, backward reasoning), which consumed a significant amount of time. The same trend can be observed in Scenario 1, though less pronounced, likely due to less time pressure and the disruption of handling the engine failure, as discussed in Section 5.1.2.
5.4. RQ4: Pilots’ Perspectives
We identified three major themes in our thematic analysis of the exit interviews and occasional comments during the think-aloud sessions.
5.4.1. Greatest added value: making a lot of information quickly accessible
For participants, the greatest added value of the DAS was not the AI features per se, but rather the quick overview of a large amount of relevant information, as explained by 30 out of 32 participants. Today, pilots must gather information from various sources, which is time-consuming and error-prone. Our DAS improves on the current state even in the Baseline version by presenting all information at a glance in the table. Several participants mentioned that the table is basically what pilots in some airlines fill out by hand as part of FOR-DEC. Having it directly available speeds up decision-making and enables better decisions by reducing workload and providing more complete information:
“If you have time, you can work it all out in half an hour. But if you have a time-critical failure, […] you need a lot of experience because you don’t have the time to work it all out. You can spend hours doing FOR-DEC, but you only have fuel for half an hour. And then you have to make an intuitive decision based on your gut feeling. These are often not the best decisions.” (C2)
But AI features do contribute to the quick overview. While having all information at a glance is helpful, it is also overwhelming, as criticized by six participants. Pilots see the importance of AI in reducing clutter by surfacing relevant and hiding or de-emphasizing less relevant information. 14 pilots stressed that recommendations and ranking by AI evaluation help them to quickly focus on good options: “I think this pre-filtering is important. And afterwards, the human can still fine-tune it a little.” (RC5). Also 14 participants explained how the color highlights help them to quickly identify potential limitations at airports: “It’s quite clever if you can see right away, do I even need to look at it or not?” (RC2). Participants also suggested how AI could further reduce clutter, e.g. by hiding very bad options, or by evaluating weather as a whole and only displaying noteworthy weather aspects. In our system, all weather components are shown in the table and evaluated individually, creating a lot of redundancy: “A little less information, visibility 10,000 or more, that’s redundant information, that doesn’t help me. So more with the question, what information creates added value?” (R5).
Two participants cautioned against careless implementations of AI. They pointed out that especially the value of recommendations may be undermined by the need to double-check them, which could slow pilots down: “You are inclined to question the computer, and you also want to decide for yourself, but at the same time you want to use everything, and then you’ll probably need ages to figure it all out.” (B8). This view is consistent with the decision times in Section 5.3, where the Rec variant turned out slower than expected.
5.4.2. Tension between more system intelligence and more bias
While pilots welcomed AI support to surface relevant information, they were also concerned that too obtrusive AI could bias them toward subpar decisions:
“The more is presented to you and processed for you, the less you are actively involved. On the one hand, it helps a lot. On the other hand, you have to be careful whether all of this is exactly what you actually want.” (C5)
“It’s a fine line as to whether you do too much.” (B1).
This tension was especially apparent with recommendations: 17 participants found recommendations helpful or suggested to add them if their system variant did not have them, while eleven participants rejected them or found them not helpful. Those who rejected recommendations felt that they remove pilots from the decision-making:
“I find it a bit difficult that the decision is given to you immediately like this [snaps fingers], and then you immediately go, ah okay, if it says so, then we’ll take Hanover.” (RC4).
“What I would not like is for the AI to simply say ‘Fly to Berlin!’ or something, because then you don’t know exactly where it’s coming from. The human is the master in the system, the AI has to be subordinate and perform supporting tasks.” (B5)
Other participants saw no threat in recommendations to their agency, emphasizing that it is their job to monitor and question the system.
Hints and highlights were seen as less problematic and accepted by all participants, as they were perceived as “less patronizing” (R5): “I have the feeling that I’m using the system, not that the system is using me.” (C2). However, seven pilots stressed that the highlights must make sense to be useful. Pilots especially disagreed with the red highlights of the “medium to good” braking action in Scenario 3, as described in Section 5.2.
Participants discussed several ways how AI may be introduced without removing pilots from the decision-making. For one, nine pilots noted that continuous support helps them to familiarize with the system and improve their SA, which also makes using the system in an emergency easier:
“It’s important to be familiar with the system, because if you only use it in an emergency, it’s like, oh yeah, is that right? If you’re always plotting where you are, you already know a bit of what it says and you don’t see, ah okay, it’s red, it doesn’t work.” (C8)
Moreover, participants had several ideas for alternative forms of AI support. Some suggested that the system only marks whether airports are suitable for a diversion or not, instead of making specific recommendations. Another idea was to display recommendations only after pilots have independently reached a decision, which some pilots likened to their collaborative decision-making:
“It’s like with our CRM [^14] Crew Resource Management, procedures for effective communication and decision-making to prevent human error., when we work together, I have to be careful not to say as captain ‘I think that airport is great, that’s where we’re flying to now, or would you mind?’ Then of course he [the first officer] says ‘No, let’s do it.’ That’s why I always have to keep it open and not say what I’ve been thinking. Let the other one say it so that there is redundancy, or perhaps have my own mistakes pointed out.” (C5)
Another suggestion by several participants was that pilots have to manually define their criteria before they get recommendations. The system could also step in with suggestions for additional criteria that the pilot may have forgotten. RC2 proposed that AI could be used to evaluate the airports according to high-level categories like weather or operations. Pilots could then sort the airports by these high-level categories, rather than by low-level criteria like in our system. C8 advocated that the system should be limited to hard facts and leave soft factors to pilots:
“The wind won’t change, that’s for sure. The fuel won’t change either. But I still have to see where the company wants us to go. Where do we perhaps know our way around? […] So these soft factors, if they were included, I think that would be a bit like taking the decision away.” (C8)
However, opinions differed on this last suggestion, as B2 argued for the opposite and said that the system should be able to judge when a situation is so critical that it is acceptable to land somewhere “even if I am now busting a limit.” (B2). The system should be able to “classify, how serious is the incident for the danger of all people on board or for the aircraft?” (B2).
While the discussions were predominantly about the risk of system-induced biases, four pilots brought the complementary perspective into the conversation, pointing out that AI could also help mitigate human errors. AI could for instance highlight good but less familiar options, and does not overlook things under time pressure.
5.4.3. Transparency and control: need for appropriation
Lastly, pilots discussed several ways in which they require transparency and control. Some participants wondered about how the AI works, e.g., why a certain recommendation is given (three participants), how the criteria are weighted (four participants), or why certain information is highlighted (four participants). However, the much more prevalent issue was what exactly the system considers for its evaluations (17 participants): Does time to destination include the time required for the approach? Does fuel at destination include final reserves? Does stop margin consider performance limitations after a technical failure? The predominant question was therefore not “How does the AI work?”, as assumed by most research in explainable AI, but rather ”How does the information fit my intention?”.
Pilots further expressed the requirement to be able to control the AI. Five participants explicitly valued the option to edit the recommendation criteria to engage with the AI in negotiation patterns, as described in Section 5.1.1. Moreover, participants made further design suggestions to enable more control, such as options to pin airports to the top of the table and to manually hide airports, or to search for airports that are not in the table. R3 further suggested to add an extra column that pilots can fill themselves, for edge cases that are not covered by the system criteria.
The common theme behind both transparency and control was that pilots want to be able to appropriate the system to fit their intention and momentary information needs. Most of the transparency and control requests were comments and questions interposed during the think-aloud sessions, where participants had an intention while using the system and asked for clarification to understand if the system would fit their intention.
6. Discussion
We discuss the takeaways of our results in Section 6.1 and Section 6.2 and bring them together in Section 6.3. We close with limitations and future work in Section 6.4.
6.1. There Is More to Decision Support than Giving Recommendations and Explanations
The common, but often tacit design goal for AI decision support is to solve the task for users, creating a redundant rather than complementary role to the human. While it is popular to speak of “human-AI collaboration” in AI-assisted decision-making research (Schmidt and Biessmann, 2020; Hemmer et al., 2023), users may perceive this recommendation-centric support more as “human-AI competition” (Zhang et al., 2023b). Some of our participants expressed similar concerns regarding recommendations.
Like previous studies involving real-world tasks with experts, our participants’ views suggest a shift in the design goal from solving the entire decision task for users to addressing their primary pain points. For the tumor assessment use case studied by Lindvall et al. (Lindvall et al., 2021), pathologists’ biggest challenge was to find small tumorous regions in huge images. In the case of sepsis diagnosis studied by Zhang et al. (Zhang et al., 2023b), physicians wanted to know which lab tests to order, when. In our case of diversions, pilots would benefit the most from a quick overview of a large amount of relevant information. This shift in design goal enables much more diverse uses for AI than merely giving end-to-end recommendations, as exemplified by the various suggestions by our participants.
The challenge shifts from giving best possible recommendations and explaining them, to serving pilots’ momentary information needs. Transparency is consequently not only—or maybe not even primarily—required to calibrate users’ reliance on the recommendation, as is the current focus in AI-assisted decision-making research (Fok and Weld, 2023). Rather, the role of transparency is to make visible how well the information fits pilots’ current intention.
Fine-grained control is further necessary to allow pilots to effectively cater the information to their intentions when the system-inferred information does not perfectly fit. This prevents the all-or-nothing situation often created by typical recommendation-centric support: Due to the closed nature of end-to-end recommendations, decision makers usually only have the choice to fully accept or reject the recommendation [^15] At least in classification tasks. In regression tasks, users can give the AI recommendation a more continuous weight in their decision-making (Poursabzi-Sangdeh et al., 2021; Green and Chen, 2019)., which is also how users’ reliance behavior is often modeled (Schemmer et al., 2023). However, as observed by Sivaraman et al. with clinicians (Sivaraman et al., 2023), decision makers’ reliance behavior can be much more nuanced when they are given the opportunity. In their case, the AI recommendation consisted of several aspects, allowing clinicians to adopt some aspect while overruling another. With our DAS, pilots could go one step further by editing the recommendation criteria. This led to some productive uses of the system in instances where the pilot would have simply ignored or rejected the system had the controls not been available.
6.2. Recommendations Have to Be Embedded into Forward Support to Be Beneficial
According to our results, recommendation-centric support significantly constrains pilots from thinking beyond the limits of the system, which in line with previous work (Jacobs et al., 2021b; Smith et al., 1997; Bussone et al., 2015) leads to more overreliance. Recommendations on their own also surprisingly did not lead to faster decisions than continuous support, as the need to review the recommendations canceled out the gains from surfacing good options. Lastly, the use of recommendations was highly disputed among participants.
But while our results suggest that recommendations may not be the most valuable use of AI for diversion assistance, we did find benefits of providing recommendations. First, recommendations can serve as confirmation, which can accelerate decisions in time-critical situations, which was particularly evident in combination with continuous support in Scenario 2. Recommendations can further benefit decision-making when they challenge pilots’ own ideas and trigger further considerations. This was more apparent in the interview statements, but could also be observed in one instance during the think-aloud sessions, namely with the participant that tried without success to align the recommendations with his own idea, as described in Section 5.1.1.
Note that both beneficial uses of recommendations require forward reasoning, as pilots must independently come up with a favored option to be confirmed or challenged. We found continuous support to be effective for promoting forward reasoning, but it lacked robustness against disruptions between normal flight and abnormal situations. Following such disruptions, pilots tended to revert to backward reasoning. To ensure reliable forward reasoning when presenting recommendations, additional measures are necessary, with potential suggestions provided by our participants.
6.3. From Recommendation-Centric to Process-Oriented Support
Bringing together the discussed aspects, we propose process-oriented decision support as a promising alternative framework to design AI decision support that is not centered around end-to-end recommendations. Figure 12 shows how process-oriented support compares to recommendation-centric support, and how it applies to the diversion use case through both continuous support and participants’ design suggestions. We consider the framework to be applicable not only for diversion assistance but more generally for AI support of high-stakes decisions. We argue that process-oriented support is particularly useful for complex decisions where human expertise and context knowledge is crucial, but hard to combine with end-to-end recommendations, such as in healthcare (Zhang et al., 2023b), social work (Kawakami et al., 2022b), or sales (Blomberg et al., 2018). The key difference to recommendation-centric support is the shift from trying to solve the task for users through end-to-end recommendations, to helping users to solve the task by addressing the challenges in their decision-making process. These challenges have to be determined for each application through user research. In our diversion use case, pilots’ main challenge was to gather the information they need to make a decision.
Figure 12. Comparison of recommendation-centric and process-oriented decision support. The yellow highlights signal the core functionality of the AI in both paradigms. The dashed arrow indicates that decision recommendations are optional in process-oriented support. The italic text lays out how process-oriented support can be implemented concretely for the diversion use case, based on our results and participants’ suggestions.
[…]
Lorem.
This shift in design goal has several implications. First, recommendation-centric support removes users from the decision-making process. It pushes users into backward reasoning, with the recommended end result as the starting point, leading to the lack of cognitive engagement and hence inappropriate reliance observed in prior work (Buçinca et al., 2021; Gajos and Mamykina, 2022). This is not only an issue for pilots as in our study, but also in other domains, such as in healthcare (Wang et al., 2019; Jacobs et al., 2021b; Fogliato et al., 2022). Process-oriented support in contrast keeps users engaged in their decision-making process and helps them to generate a decision themselves while reasoning forward. Recommendations are optional, serving as confirmation or challenge toward the end of the process rather than the starting point. As reflected by our participants’ discussions, the core challenge for process-oriented support is to strike a good balance between interpreting information with AI and leaving the interpretation work to the human, which is strongly use-case-dependent. Our continuous support concept leaves much of the interpretation to pilots, which worked well for the diversion use case since pilots are highly trained expert users who usually know what information is important, as demonstrated by participants in the Baseline condition. For such expert users, a restrained form of AI support like continuous support has the benefit that it allows them to work very flexibly with the system, as discussed by our participants (Section 5.4.2). Use cases with less expert users may require a “stronger” role for the AI, e.g., guiding users toward important information through explanations. Even for our use case, some pilots wished for more sophisticated AI functionality.
Second, the role of transparency changes. In recommendation-centric support, transparency aids appropriate reliance on recommendations. In process-oriented support, transparency facilitates appropriation (Dix, 2007) by helping users understand how well the system functionality aligns with their momentary intentions and needs. The importance of appropriation in AI-assisted decision-making has been discussed by several authors recently (Zhang et al., 2023a; Cabitza et al., 2021; Ehsan et al., 2021). It differs from reliance calibration in at least two aspects. One is that transparency for appropriation is targeting the more granular level of intermediary decision-making steps, rather than the end of the process. This leads to more nuanced behavior than a simple accept-reject dichotomy, like partial reliance and negotiation patterns, as observed by Sivaraman et al. (Sivaraman et al., 2023) and with our participants as well. The other difference is that appropriate reliance concerns the correctness of the AI, while appropriation is about the user’s intention. As was evident from our participants’ requests, this does not necessarily involve explaining how the algorithm works.
Lastly, while control has been studied in AI systems for instance in the context of interactive machine learning (Kulesza et al., 2015; Dudley and Kristensson, 2018; Smith-Renner et al., 2020), there is usually no control in recommendation-centric decision support. In process-oriented support, control is an important complement to transparency to enable appropriation. Notably, different to interactive machine learning, control is not necessarily about feedback that the model should learn, but rather about steering the system according to users’ momentary intentions and context, as exemplified by the criteria editing feature in our system, or the controls designed by Cai et al. in a healthcare application (Cai et al., 2019).
Process-oriented support joins the ranks of recent alternatives to recommendation-centered support. In particular, we see it as a generalization of evaluative AI (Miller, 2023). Core to process-oriented support is to identify the main challenges in a decision-making process which can be supported by AI. Evaluative AI can be interpreted as process-oriented support for decisions where the main challenge is to generate and evaluate multiple hypotheses, such as in medical diagnostics. In the diversion use case, the main challenge was different, namely the difficulty of information gathering. This leads to different support opportunities, such as our continuous support concept, which proved to be a good, but not perfect starting point. Further research is required to better understand how each element of process-oriented support can be implemented for diversion assistance and beyond, from useful AI support roles, over how to enable appropriation through transparency and control, to how recommendations can be integrated toward the end of users’ decision-making process.
6.4. Limitations and Future Work
There are several limitations to our results, mostly stemming from our flight simulation setup. For one, we decided to conduct the study with a single pilot at a time, rather than with two pilots as is the standard in today’s cockpits. We are fully aware that this is a significant deviation from how pilots work today, as the cooperation of both pilots is fundamental to today’s operations (Helmreich et al., 1999). We still decided for a single-pilot setup for two reasons. First, recruiting pilots is challenging, and relying on the simultaneous availability of pairs of pilots would have significantly reduced our sample size or even made the study impossible to conduct. Second, we are interested in how different AI support paradigms would affect pilots’ decision-making in a possible AI-augmented future of aviation. We are agnostic as to whether this future maintains a two-pilots cockpit or involves single-pilot operations, which is a long-standing goal of the aviation industry (Harris, 2007).
Another limitation was that we could not provide all tools that pilots use in their daily work, such as maps or tools to calculate certain performance and flight data. Besides being prohibitively costly to replicate all these tools, they are also not standardized between airlines. Still, participants were able to appropriate what was available to them in X-Plane to fulfill many of their information needs. Several participants also praised the study setup to be well executed. We therefore assume that the impact of the limited tools was not critical.
Furthermore, despite our best efforts to construct a representative range of valid scenarios, it was not possible in our setup to fully reflect complexities that can arise from the coordination with multiple stakeholders. Such coordination was rudimentarily present in Scenario 3, but in reality, these interactions would be much more open-ended and therefore hard to simulate in a controlled setting like ours. While studies like ours do produce valuable insights, other methods would be required to better capture the intricacies of such open-ended scenarios. For instance, one could analyze incident reports post deployment, which is a proven strategy in aviation to improve safety.
Lastly, given the difficulty and cost to recruit professional pilots, our sample size was relatively small for our quantitative analyses. We tried to address this limitation by triangulating the quantitative results with rich qualitative data from the think-aloud sessions and exit interviews. We found clear trends in the quantitative data that were consistent with the qualitative data.
Overall, these limitations highlight the difficulty of rigorously evaluating different support paradigms on a realistic task with domain experts. We suspect this is also the reason why such studies are rare. Nevertheless, similar to previous AI-assisted decision-making research on real-world tasks, our study produced insights that would likely not be available with simple tasks and crowd worker participants. We therefore encourage further research on realistic tasks with domain experts to close the gap between research and real-world adoption of AI decision support.
7. Conclusion
We conducted one of the first empirical comparisons between the dominant recommendation-centric AI decision support paradigm and alternative approaches in a realistic task setting with domain experts. We found that with recommendation-only, participants exhibited strong overreliance as they thought beyond the limits of the system significantly less than with the baseline variant. The recommendation-only variant was also surprisingly not faster than continuous support, in fact it even tended to be slower. We further found that to benefit from recommendations, pilots must engage with them while reasoning forward. Continuous support appeared to be an effective approach to encourage forward reasoning, allowing to display recommendations with less overreliance and leading to faster decisions when combined with recommendations. However, the effectiveness of continuous support was sensitive to disruptions between normal flight and abnormal situations. Our qualitative analysis reveals that pilots do not primarily value the system for suggesting suitable options, but rather for helping them to gain a quick overview of a large amount of information. While pilots welcome AI features for this purpose, they are concerned that too obtrusive AI takes away too much of the decision from them, with recommendations being particularly controversial. Participants’ statements further revealed their requirements for transparency and control to appropriate the system. Notably, their primary concern with transparency was not how the AI works, but rather how it fits their intention.
Our results challenge the assumption that AI decision support should be recommendation-centric and highlight the importance of supporting the decision-making process in a forward direction. Our continuous support concept was a promising first step in this direction, while our findings suggest many opportunities for improvement and future work. Further research is especially required to understand how to more robustly support decisions in a forward direction. We encapsulated our findings in a framework for process-oriented support, and envision our work as a contribution toward a more holistic perspective on AI-assisted decision-making that looks beyond recommendations and explanations.
Acknowledgements.
This work was supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) under the LuFo VI-1 program, project KIEZ4-0. Many thanks to our participants, who were eager to discuss our research with us, were happy to provide us with valuable contacts and insights into their daily work, and spread the word about our study among their colleagues.
Appendix A Exit Interview Guide
  • How was your impression of the system?
  • On a scale from 1—not helpful at all to 5—extremely helpful, how do you rate the helpfulness of the system?
  • What is the greatest added value of the system?
  • What do you find problematic or in need of improvement about the system?
  • What was your strategy for using the system? What did you use it for?
  • For Rec, Cont, Rec+Cont conditions: How important was/were the recommendations/normal flight mode/color highlights for your usage?
  • For Baseline condition: What kind of AI features would you like to see in the system?
  • Anything else that you want to ask or comment on?
Appendix B Participant Details
Table 3. Focus group participant details.
ID Gender Rank Flight hours Participated in pilot test
F1 Male F/O 400
F2 Male Captain 10 2001020010\,20010 200
F3 Male F/O 1600 X
F4 Male F/O 370 X
Median (IQR) 1000 (392.5–3750)
Table 4. Study participant details.
ID Gender Rank Flight hours # Diversions
B1 Male F/O 3180 0
B2 Male F/O 4000 4
B3 Male F/O 2000 4
B4 Male F/O 2000 1
B5 Male F/O 200 0
B6 Male F/O 1700 2
B7 Male F/O 4800 3
B8 Male F/O 5000 1
R1 Male F/O 4000 2
R2 Male F/O 2000 1
R3 Male F/O 2500 1
R4 Male F/O 4000 5
R5 Male F/O 3000 0
R6 Male F/O 1640 0
R7 Male Captain 11 0001100011\,00011 000 1
R8 Male F/O 2500 1
C1 Male F/O 2400 0
C2 Male F/O 3000 0
C3 Male F/O 6000 7
C4 Male F/O 2500 2
C5 Male Captain 13 0001300013\,00013 000 5
C6 Male F/O 2500 0
C7 Male F/O 4000 5
C8 Female F/O 580 0
RC1 Male F/O 3800 3
RC2 Male F/O 4000 0
RC3 Male F/O 2000 3
RC4 Male F/O 2400 2
RC5 Male F/O 1900 3
RC6 Male F/O 800 0
RC7 Female F/O 2500 2
RC8 Male F/O 3000 2
Median (IQR) 2500 (2000–4000) 1.5 (0–3)